The length of a rectangle is 3 cm more than 5 times its width. If the area of the rectangle is 76 cm2, find the dimensions of the rectangle to the nearest thousandth
I started with x^2+(x+3)^+5=76, but it does not make sense.

Answer by scott8148(6628)

let x=width; so 5x+3=length ... area=(length)(width)

76=(5x+3)(x) ... 76=5x^2+3x ... 0=5x^2+3x-76

use quadratic formula to find x